OpenStudy (anonymous):

WILL FAN AND MEDAL! need help making a problem Sheldon and Leonard are preparing for their annual trip to Comic-Con, a comic book convention. After careful analysis of route options, and based on traffic patterns, Leonard determines the following: while driving at a constant pace, they will travel 50 miles in the first hour, 255 miles in the next 3 hours, and 120 miles in the next 2 hours. Create a piecewise function representing Leonard and Sheldon’s traveling time and distance.

OpenStudy (anonymous):

if distance =y and time=x

OpenStudy (anonymous):

y is distance and x is time?

OpenStudy (anonymous):

yes! @satellite73

OpenStudy (anonymous):

ok then for \(0\leq x\leq 1\) you know \(y=50x\) as distance equals rate times time

OpenStudy (anonymous):

what is the rate for the next three hours?

OpenStudy (anonymous):

255?

OpenStudy (anonymous):

on no that is the total distance for 3 hours to find the rate, divide 255 by 3

OpenStudy (anonymous):

85

OpenStudy (anonymous):

wow they are going fast

OpenStudy (anonymous):

so so far y=50x+85...?

OpenStudy (anonymous):

no you are looking for a piecewise function right?

OpenStudy (anonymous):

yess..

OpenStudy (anonymous):

going to look something like \[y = \left\{\begin{array}{rcc} 50x & \text{if} & 0\leq x\leq 1\\ 50+85x& \text{if} &1<x\leq 4 \end{array} \right.\]

OpenStudy (anonymous):

what is the next rate?

OpenStudy (anonymous):

120 miles in two hours?

OpenStudy (anonymous):

thats divided by 3?

OpenStudy (anonymous):

is what divided by three?

OpenStudy (anonymous):

would that be 40?

OpenStudy (anonymous):

120 miles in two hours?

OpenStudy (anonymous):

oh scuse me sorry, 60?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok now only one more step how far have they gone so far?

OpenStudy (anonymous):

for the first four hours that is

OpenStudy (anonymous):

425 miles

OpenStudy (anonymous):

in all?

OpenStudy (anonymous):

no for the first four hours

OpenStudy (anonymous):

well for the first hr it says 50

OpenStudy (anonymous):

yes and then how far in the next three hours?

OpenStudy (anonymous):

305?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so its the equation you wrote down and =305?

OpenStudy (anonymous):

\[y = \left\{\begin{array}{rcc} 50x & \text{if} & 0\leq x\leq 1\\ 50+85x& \text{if} &1<x\leq 4\\ 305+60x&\text{ if} &4<x\leq6 \end{array} \right.\]

OpenStudy (anonymous):

thats the piecewise function?